what is the atomic spacing between two iron atoms?


1
Expert's answer
2021-04-06T06:31:54-0400

To be given in question

Atom= iron atom


To be asked in question

Atomic space in iron atom


We know that

Room temperature iron atom is bcc type structure



BCC type structure


Use pithagorse theorem

"(bd)^2=a^2+(fd)^2"

"(bd)^2=a^2+(2a^2)"

"(4r)^2=3a^2"

"a=\frac{4r}{√3}"

Inter spacing between iron atom


"2r=\frac{a√3}{2}"